# Classical harmonic analysis viewed through the prism of noncommutative geometry

### v4

@@ 1 Introduction
If $E$ is a Hermitian complex vector bundle over a closed smooth manifold $M$ endowed with a Lebesgue measure, it is interesting to observe that if $D\colon\mathscr{C}^{\infty}(M,E)\to\mathscr{C}^{\infty}(M,E)$ is a Hermitian elliptic pseudo-differential operator of order $m>0$ acting on the space $\mathscr{C}^{\infty}(M,E)$ of smooth sections of $E$ then it induces a selfadjoint operator on the Hilbert space $\mathrm{L}^{2}(M,E)$. Moreover, the operator $\operatorname{\mathrm{sign}}D$ induces a pseudo-differential operator of order 0 acting on $\mathscr{C}^{\infty}(M,E)$, see [BaG82b, Proposition 2.4, p. 143] and its proof. Recall that the multiplication operator $\pi(f)\colon\mathrm{L}^{2}(M,E)\to\mathrm{L}^{2}(M,E)$, defined by any function $f\in\mathrm{C}(M)$ acting on the Hilbert space, satisfies a standard property: the commutator $P,\pi(f)]$ of any pseudo-differential operator $P$ of order 0 is a compact operator. This entails that any Hermitian elliptic pseudo-differential operator defines a K-homology class $\operatorname{\mathrm{sign}}D]$, often denoted $[D]$, in the K-homology group $\mathrm{K}^{1}(\mathrm{C}(M))$. Using the operator $\operatorname{\mathrm{sign}}\begin{bmatrix}0&D^{*}\\ D&0\\ \end{bmatrix}$ for an elliptic pseudo-differential operator $D$ of order $m>0$, instead of $\operatorname{\mathrm{sign}}D$, we can also obtain a K-homology class $[D]$ in the even K-homology group $\mathrm{K}^{0}(\mathrm{C}(M))$.
@@
by integrating over the boundary of a larger bisector $\Sigma^{\mathrm{bi}}_{\nu}$ using the resolvent operator $R(z,D)\overset{\mathrm{def}}{=}(z-D)^{-1}$, where $\omega<\nu<\theta$. Note that the boundedness of such a functional calculus is not free, contrary to the case of the functional calculus of a selfadjoint operator. We refer to our paper [Arh24c] for concrete examples of operator with such a bounded functional calculus, where we use a notion of curvature for obtaining it. Using the function $\operatorname{\mathrm{sgn}}$ defined by $\operatorname{\mathrm{sgn}}(z)\overset{\mathrm{def}}{=}1_{\Sigma_{\theta}}(z)-1_{-\Sigma_{\theta}}(z)$, we show as the hilbertian case that $\operatorname{\mathrm{sign}}D\colon X\to X$ is a Fredholm module. This notion of functional calculus was popularized in the paper [AKM06], which contains a (second) solution to famous Kato’s square root problem solved in [AHLMT02] and in [AKM06] (see also [HLM02] and [Tch01] and the Bourbaki seminar [Mey03]).
@@
(1.11) $\operatorname{\mathrm{sign}}D=D(D^{2})^{-\frac{1}{2}}=\begin{bmatrix}0&*\\ \frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}(-\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}M_{a}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x})^{-\frac{1}{2}}&0\end{bmatrix}$.
@@
(1.14) $\operatorname{\mathrm{sign}}D=D(D^{2})^{-\frac{1}{2}}=\begin{bmatrix}0&*\\ \partial A^{-\frac{1}{2}}&0\end{bmatrix}$.
From this, it is apparent that the vectorial Riesz transform $\partial A^{-\frac{1}{2}}\colon\mathrm{L}^{p}(\Omega)\to\mathcal{X}_{p}$ is bounded and appears in the operator $F\overset{\mathrm{def}}{=}\operatorname{\mathrm{sign}}D$ and particularly with the possible pairing with the group $\mathrm{K}_{0}(\mathcal{A})$ of $\mathrm{K}$-theory for a suitable subalgebra $\mathcal{A}$ of the algebra $\mathrm{L}^{\infty}(\Omega)$, as we will see. Of course, we can consider more generally sub-Markovian semigroups acting on the noncommutative $\mathrm{L}^{p}$-space $\mathrm{L}^{p}(\mathcal{M})$ of a von Neumann algebra $\mathcal{M}$ in the previous discussion.
@@
This paper is structured as follows. Section 2 provides the necessary background and revisits key notations, as well as essential results required for our work. In Section 3, we introduce the concept of (odd or even) Banach Fredholm modules and define new groups, $\mathrm{K}^{0}(\mathcal{A},\mathscr{B})$ and $\mathrm{K}^{1}(\mathcal{A},\mathscr{B})$, in $\mathrm{K}$-homology, associated with an algebra $\mathcal{A}$ and a suitable class $\mathscr{B}$ of Banach spaces. Section 4 focuses on establishing pairings with the $\mathrm{K}$-theory groups of the algebra $\mathcal{A}$, which is central to our approach. In Section 5, we introduce a suitable notion of summability for Banach Fredholm modules, revisiting key concepts related to approximation numbers of operators acting on Banach spaces. Section 6 is dedicated to the Chern character associated with finitely summable Banach Fredholm modules. Following the classical Hilbert space approach, we demonstrate how the pairing between $\mathrm{K}$-homology and $\mathrm{K}$-theory groups can be described using the Chern character. In Section 7, we revisit the notion of Banach spectral triples and explain how Banach Fredholm modules can be constructed from these triples. Section 8 explores concrete examples of triples constructed using Dirac operators, leading to Banach Fredholm modules, where we compute the pairings with $\mathrm{K}$-theory in specific cases. We also reveal the connection between the summability of Banach spectral triples and Sobolev embedding theorems. In Section 9, we present examples of Banach Fredholm modules in the context of noncommutative Hardy spaces, leading to new Banach Fredholm modules on group $\mathrm{C}^{*}$-algebras, even within the Hilbert space framework. Section 10 present a Banach Fredholm module on the reduced $\mathrm{C}^{*}$-algebra of the free group $\mathbb{F}_{\infty}$. Finally, in Section 11 discusses how to associate a Banach spectral triple with each reasonable sub-Markovian semigroup. We also observe that the corresponding Banach Fredholm module, which may be kernel-degenerate, is closely linked to the vectorial Riesz transform associated to the semigroup.
@@ 2 Preliminaries
(invariance under compact perturbations). If $T\colon X\to Y$ and $S\colon Y\to Z$ are Fredholm operators then the composition $ST$ is also a Fredholm operator and $\operatorname{Index}ST=\operatorname{Index}S+\operatorname{Index}T$, see [AbA02, Theorem 4.43 p. 158]. Moreover, by [AbA02, Theorem 4.42 p. 157], a bounded operator $T\colon X\to Y$ is a Fredholm operator if and only if its adjoint $T^{*}\colon X^{*}\to Y^{*}$ is a Fredholm operator. In this case, we have $\operatorname{Index}T^{*}=-\operatorname{Index}T$.
@@
With the function $\operatorname{\mathrm{sgn}}\overset{\mathrm{def}}{=}1_{\Sigma_{\theta}}-1_{-\Sigma_{\theta}}$, we can define the bounded operator $\operatorname{\mathrm{sgn}}(A)\colon X\to X$.
@@ 3 Banach Fredholm modules and Banach $\mathrm{K}$-homology
For any $1\leqslant p<\infty$, it is natural to consider the class $\mathscr{L}^{p}$ of spaces isomorphic to $\mathrm{L}^{p}$-spaces and the class $\mathscr{L}^{p}_{\mathrm{nc}}$ of spaces isomorphic to noncommutative $\mathrm{L}^{p}$-spaces. We could also consider the class $\mathscr{S}\mathscr{L}_{p}$ of saces isomorphoc to subspaces of $\mathrm{L}^{p}$-spaces and the class $\mathscr{S}\mathscr{Q}\mathscr{L}_{p}$ of spaces isomorphic to subspaces of quotients of $\mathrm{L}^{p}$-spaces.
@@ 4 Coupling between Banach K-homology with K-theory
While Lafforgue’s $\mathrm{K}\mathrm{K}^{\mathrm{ban}}$ provides a powerful bivariant framework and an abstract product formalism [Laf02], it does not supply a concrete Toeplitz-type formula for the pairing. The novelty here is precisely this explicit, calculable index formula in a non-Hilbert Banach context. The construction applies uniformly to classes $\mathscr{B}$ stable finite sums (e.g., the class of $\mathrm{L}^{p}$-spaces up to isomorphism), and it will serve as the analytic backbone for the Banach-space Chern characters developed later in the paper.
@@ 5 Summability of Banach Fredholm modules
The Weyl numbers are defined by
@@
We refer to [Pie07, Section 6.2] for more examples of $s$-number sequences.
@@ 6 The Chern character of a $q$-summable Banach Fredhlom module
$=P([P,u^{-1}])uP=\frac{1}{2}P\mathop{}\mathopen{}\mathrm{d}u^{-1}uP$
$=\frac{1}{2}P\mathop{}\mathopen{}\mathrm{d}u^{-1}(uP-Pu+Pu)$
@@
Consequently, this operator also belongs to the space $S^{N}_{\mathrm{app}}(X)$. So we use Corollary 6.5 with the operators $PuP\colon P(X)\to P(X)$ and $Pu^{-1}P\colon P(X)\to P(X)$ replacing $T$ and $R$. Using the equalities
$\operatorname{Tr}(\mathop{}\mathopen{}\mathrm{d}u^{-1}\mathop{}\mathopen{}\mathrm{d}u)^{N}=\operatorname{Tr}(\mathop{}\mathopen{}\mathrm{d}u\mathop{}\mathopen{}\mathrm{d}u^{-1})^{N}$
and
@@ 7 From Banach spectral triples to Banach Fredhlom modules
If the second condition is replaced by the weaker assumption $D^{-1}$ is a compact operator on $\overline{\operatorname{Ran}D}$¿¿, the corresponding structure is referred to as a possibly kernel-degenerate compact Banach spectral triple.
@@ 8 Examples of triples and Banach Fredholm modules
It can also be viewed as a Fourier multiplier with symbol $\mathrm{i}\,\operatorname{\mathrm{sign}}(n)$, i.e., we have
@@
Second proof of the first assertion. Observe that by $|D|$ is a pseudo-differential operator of order $1$. Moreover, recall that if $f\in\mathrm{C}^{\infty}(\mathbb{T})$ then $M_{f}$ is a differential operator of order 0, hence a pseudo-differential operator of order 0 by [Dieu88, Remark 23.26.5]. Consequently, according to [RuT10, Remark 5.3.3 p. 423] the commutator $|D|,M_{f}]$ is a pseudo-differential operator of order 0. Such operator induces a bounded operator on the Banach space $\mathrm{L}^{p}(\mathbb{T})$ by [RuT10, Theorem 5.2.23 p. 420].

### v5

@@ 1 Introduction
If $E$ is a Hermitian complex vector bundle over a closed smooth manifold $M$ endowed with a Lebesgue measure, it is interesting to observe that if $D\colon\mathscr{C}^{\infty}(M,E)\to\mathscr{C}^{\infty}(M,E)$ is a Hermitian elliptic pseudo-differential operator of order $m>0$ acting on the space $\mathscr{C}^{\infty}(M,E)$ of smooth sections of $E$ then it induces a selfadjoint operator on the Hilbert space $\mathrm{L}^{2}(M,E)$. Moreover, the operator $\mathrm{sign}D$ induces a pseudo-differential operator of order 0 acting on $\mathscr{C}^{\infty}(M,E)$, see [BaG82b, Proposition 2.4, p. 143] and its proof. Recall that the multiplication operator $\pi(f)\colon\mathrm{L}^{2}(M,E)\to\mathrm{L}^{2}(M,E)$, defined by any function $f\in\mathrm{C}(M)$ acting on the Hilbert space, satisfies a standard property: the commutator $P,\pi(f)]$ of any pseudo-differential operator $P$ of order 0 is a compact operator. This entails that any Hermitian elliptic pseudo-differential operator defines a K-homology class $\mathrm{sign}D]$, often denoted $[D]$, in the K-homology group $\mathrm{K}^{1}(\mathrm{C}(M))$. Using the operator $\mathrm{sign}\begin{bmatrix}0&D^{*}\\ D&0\\ \end{bmatrix}$ for an elliptic pseudo-differential operator $D$ of order $m>0$, instead of $\mathrm{sign}D$, we can also obtain a K-homology class $[D]$ in the even K-homology group $\mathrm{K}^{0}(\mathrm{C}(M))$.
@@
by integrating over the boundary of a larger bisector $\Sigma^{\mathrm{bi}}_{\nu}$ using the resolvent operator $R(z,D)\overset{\mathrm{def}}{=}(z-D)^{-1}$, where $\omega<\nu<\theta$. Note that the boundedness of such a functional calculus is not free, contrary to the case of the functional calculus of a selfadjoint operator. We refer to our paper [Arh24c] for concrete examples of operator with such a bounded functional calculus, where we use a notion of curvature for obtaining it. Using the function $\operatorname{\mathrm{sgn}}$ defined by $\operatorname{\mathrm{sgn}}(z)\overset{\mathrm{def}}{=}1_{\Sigma_{\theta}}(z)-1_{-\Sigma_{\theta}}(z)$, we show as the hilbertian case that $\mathrm{sign}D\colon X\to X$ is a Fredholm module. This notion of functional calculus was popularized in the paper [AKM06], which contains a (second) solution to famous Kato’s square root problem solved in [AHLMT02] and in [AKM06] (see also [HLM02] and [Tch01] and the Bourbaki seminar [Mey03]).
@@
(1.11) $\mathrm{sign}D=D(D^{2})^{-\frac{1}{2}}=\begin{bmatrix}0&*\\ \frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}(-\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}M_{a}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x})^{-\frac{1}{2}}&0\end{bmatrix}$.
@@
(1.14) $\mathrm{sign}D=D(D^{2})^{-\frac{1}{2}}=\begin{bmatrix}0&*\\ \partial A^{-\frac{1}{2}}&0\end{bmatrix}$.
From this, it is apparent that the vectorial Riesz transform $\partial A^{-\frac{1}{2}}\colon\mathrm{L}^{p}(\Omega)\to\mathcal{X}_{p}$ is bounded and appears in the operator $F\overset{\mathrm{def}}{=}\mathrm{sign}D$ and particularly with the possible pairing with the group $\mathrm{K}_{0}(\mathcal{A})$ of $\mathrm{K}$-theory for a suitable subalgebra $\mathcal{A}$ of the algebra $\mathrm{L}^{\infty}(\Omega)$, as we will see. Of course, we can consider more generally sub-Markovian semigroups acting on the noncommutative $\mathrm{L}^{p}$-space $\mathrm{L}^{p}(\mathcal{M})$ of a von Neumann algebra $\mathcal{M}$ in the previous discussion.
@@
This paper is structured as follows. Section 2 provides the necessary background and revisits key notations, as well as essential results required for our work. In Section 3, we introduce the concept of (odd or even) Banach Fredholm modules and define new groups, $\mathrm{K}^{0}(\mathcal{A},\mathscr{B})$ and $\mathrm{K}^{1}(\mathcal{A},\mathscr{B})$, in $\mathrm{K}$-homology, associated with an algebra $\mathcal{A}$ and a suitable class $\mathscr{B}$ of Banach spaces. Section 4 focuses on establishing pairings with the $\mathrm{K}$-theory groups of the algebra $\mathcal{A}$, which is central to our approach. In Section 5, we introduce a suitable notion of summability for Banach Fredholm modules, revisiting key concepts related to approximation numbers of operators acting on Banach spaces. Section 6 is dedicated to the Chern character associated with finitely summable Banach Fredholm modules. Following the classical Hilbert space approach, we demonstrate how the pairing between $\mathrm{K}$-homology and $\mathrm{K}$-theory groups can be described using the Chern character. In Section 7, we revisit the notion of Banach spectral triples and explain how Banach Fredholm modules can be constructed from these triples. Section 8 explores concrete examples of triples constructed using Dirac operators, leading to Banach Fredholm modules, where we compute the pairings with $\mathrm{K}$-theory in specific cases. We also reveal the connection between the summability of Banach spectral triples and Sobolev embedding theorems. In Section 9, we present examples of Banach Fredholm modules in the context of noncommutative Hardy spaces, leading to new Banach Fredholm modules on group $\mathrm{C}^{*}$-algebras, even within the Hilbert space framework. Section 9.5 present a Banach Fredholm module on the reduced $\mathrm{C}^{*}$-algebra of the free group $\mathbb{F}_{\infty}$. Finally, in Section 10 discusses how to associate a Banach spectral triple with each reasonable sub-Markovian semigroup. We also observe that the corresponding Banach Fredholm module, which may be kernel-degenerate, is closely linked to the vectorial Riesz transform associated to the semigroup.
@@ 2 Preliminaries
(invariance under compact perturbations). If $T\colon X\to Y$ and $S\colon Y\to Z$ are Fredholm operators then the composition $ST$ is also a Fredholm operator and $\operatorname{Index}ST=\operatorname{Index}S+\operatorname{Index}T$, see [AbA02, Theorem 4.43 p. 158]. Moreover, by [AbA02, Theorem 4.42 p. 157], a bounded operator $T\colon X\to Y$ is a Fredholm operator if and only if its adjoint $T^{*}\colon X^{*}\to Y^{*}$ is a Fredholm operator. In this case, we have
(2.3) $\operatorname{Index}T^{*}=-\operatorname{Index}T.$
@@
With the function $\operatorname{\mathrm{sgn}}\overset{\mathrm{def}}{=}1_{\Sigma_{\theta}}-1_{-\Sigma_{\theta}}$, we can define the bounded operator $\operatorname{\mathrm{sgn}}D\colon X\to X$.
@@ 3 Banach Fredholm modules and Banach $\mathrm{K}$-homology
For any $1\leqslant p<\infty$, it is natural to consider the class $\mathscr{L}^{p}$ of spaces isomorphic to $\mathrm{L}^{p}$-spaces and the class $\mathscr{L}^{p}_{\mathrm{nc}}$ of spaces isomorphic to noncommutative $\mathrm{L}^{p}$-spaces. We could also consider the class $\mathscr{S}\mathscr{L}_{p}$ of saces isomorphoc to subspaces of $\mathrm{L}^{p}$-spaces and the class $\mathscr{S}\mathscr{Q}\mathscr{L}_{p}$ of spaces isomorphic to subspaces of quotients of $\mathrm{L}^{p}$-spaces. Finally, we can use the class $\mathrm{Ban}$ of all Banach spaces.
@@ 4 Coupling between Banach K-homology with K-theory
While Lafforgue’s $\mathrm{K}\mathrm{K}^{\mathrm{ban}}$ provides a powerful bivariant framework and an abstract product formalism [Laf02], it does not supply a concrete Toeplitz-type formula for the pairing. The novelty here is precisely this explicit, calculable index formula in a non-Hilbert Banach context. The construction applies uniformly to classes $\mathscr{B}$ stable finite sums (e.g., the class $\mathscr{L}^{p}$ of $\mathrm{L}^{p}$-spaces up to isomorphism or the class $\mathrm{Ban}$ of all Banach spaces), and it will serve as the analytic backbone for the Banach-space Chern characters developed later in the paper.
@@ 5 Summability of Banach Fredholm modules
These numbers quantify how well the map can be approximated by finite-dimensional linear operators. The Weyl numbers are defined by
@@
We refer to [Pie07, Section 6.2] for more examples of $s$-number sequences. In Section 8.2, we will use the following observation.
Lemma 5.2
Suppose that $1\leqslant p\leqslant\infty$. Consider two $\ell^{p}$-direct sums $X\overset{\mathrm{def}}{=}\bigoplus_{k=1}^{N}X_{k}$ and $Y\overset{\mathrm{def}}{=}\bigoplus_{k=1}^{N}Y_{k}$ of Banach spaces and a block diagonal operator $T\overset{\mathrm{def}}{=}\bigoplus_{k=1}^{N}T_{k}\colon X\to Y$, where each $T_{k}\colon X_{k}\to Y_{k}$ is a bounded linear operator. Then for every integer $n\geqslant 1$ we have
(5.6) $a_{Nn}(T)\leqslant\max_{1\leqslant k\leqslant N}a_{n}(T_{k}).$
Proof : Let $n\geqslant 1$ and $\varepsilon>0$. For each $k\in\{1,\dots,N\}$ choose a finite-rank operator $S_{k}\colon X_{k}\to Y_{k}$ with $\operatorname{rank}S_{k}<n$ and $\left\|T_{k}-S_{k}\right\|\leqslant a_{n}(T_{k})+\varepsilon$. Consider the block diagonal operator $S\overset{\mathrm{def}}{=}\bigoplus_{k=1}^{N}S_{k}\colon X\to Y$. Then $\operatorname{rank}S\leqslant\sum_{k=1}^{N}\operatorname{rank}S_{k}<Nn$. Moreover, by the definition of the operator norm on the direct sum,
$\left\|T-S\right\|=\sup_{1\leqslant k\leqslant N}\left\|T_{k}-S_{k}\right\|\leqslant\sup_{1\leqslant k\leqslant N}\big(a_{n}(T_{k})+\varepsilon\big)=\max_{1\leqslant k\leqslant N}a_{n}(T_{k})+\varepsilon.$
Taking the infimum over all such families $(S_{k})_{1\leqslant k\leqslant N}$ and letting $\varepsilon\to 0$ yields (5.6).
@@ 6 The Chern character of a $q$-summable Banach Fredhlom module
$=P([P,u^{-1}])uP=\frac{1}{2}P\mathop{}\mathopen{}\mathrm{d}u^{-1}uP=\frac{1}{2}P\mathop{}\mathopen{}\mathrm{d}u^{-1}(uP-Pu+Pu)$
@@ 7 From Banach spectral triples to Banach Fredhlom modules
If the second condition is replaced by the weaker assumption $D^{-1}$ is a compact operator on $\overline{\operatorname{Ran}D}$¿¿, the corresponding structure is referred to as a possibly kernel-degenerate compact Banach spectral triple. If in addition, there exists an isomorphism $\gamma\colon X\to X$, with $\gamma^{2}=\mathrm{Id}$,
(7.2) $D\gamma=-\gamma D\quad\text{and}\quad\gamma\pi(a)=\pi(a)\gamma,\quad a\in\mathcal{A},$
the spectral triple is said to be even or graded. Otherwise, it is said to be odd or ungraded.
@@ 8 Examples of triples and Banach Fredholm modules
We describe in details some examples. We also refer to [Arh26] for the case of Hodge-Dirac operators on oriented compact Riemannian manifolds.
@@
It can also be viewed as a Fourier multiplier with symbol $\mathrm{i}\,\mathrm{sign}(n)$, i.e., we have
@@
Second proof of the first assertion. Observe that $|D|$ is a pseudo-differential operator of order $1$. Moreover, recall that if $f\in\mathrm{C}^{\infty}(\mathbb{T})$ then $M_{f}$ is a differential operator of order 0, hence a pseudo-differential operator of order 0 by [Dieu88, Remark 23.26.5]. Consequently, according to [RuT10, Remark 5.3.3 p. 423] the commutator $|D|,M_{f}]$ is a pseudo-differential operator of order 0. Such operator induces a bounded operator on the Banach space $\mathrm{L}^{p}(\mathbb{T})$ by [RuT10, Theorem 5.2.23 p. 420].